OpenStudy (zyberg):

If n and 6 are coprime numbers, prove that n^2 - 1 is divisible by 24.

OpenStudy (zyberg):

I got just as far as to write n^2 - 1 = 24 * k

jhonyy9 (jhonyy9):

in trhis case n=6+/-1 so n=5 and/or 7 for 5 will get 25-1=24 what is divisidle by 24 for 7 ----- 49-1=48=24*2 ------- 24

OpenStudy (zyberg):

@jhonyy9 I need more universal proof. :/

jhonyy9 (jhonyy9):

i thought it on complet induction proof method 6=3*2 5=3*2-1 7=3*2+1 use for this 6n and 6n +/- 1

jhonyy9 (jhonyy9):

for 5 and 7 this is true like above wrote so suppose for n=k is true and rewrite it and proove it for (k+1)

jhonyy9 (jhonyy9):

for 6n will be n^2 -1 so (6n)^2 -1 = (6n-1)(6n+1) for case of 6n-1 will result (6n-1)^2 -1 = (6n)^2 -12n +1 and for 6n+1 ------- (6n+1)^2 -1 = (6n)^2 +12n +1

ganeshie8 (ganeshie8):

Step1 Consider the remainders when 'n' is divided by 6

OpenStudy (zyberg):

5,4,3,2,1

ganeshie8 (ganeshie8):

What remainders are possible ? Keep in mind the given fact : n and 6 are coprime

OpenStudy (zyberg):

So, the possible remainders are 5, 4, 3, 2, and 1.

ganeshie8 (ganeshie8):

Really sure the remainders 2,3,4 are valid ?

OpenStudy (zyberg):

Oh, you are right! So, they are 5, and 1.

ganeshie8 (ganeshie8):

Why ? I'd like to see your explanation for excluding 2,3,4,0

OpenStudy (zyberg):

So, 6 can be divided by 2, 3. To get remainder 2, we need to have number of form 6k + 2; n would need to be 6k+2, but n and 6 are coprime and 6k+2 = 2(3k +1) Same goes for 3 and 0. Not really sure about 4 :/

ganeshie8 (ganeshie8):

Nice. Is it hard to see that n and 6 are not coprime when n = 6k+4 ?

OpenStudy (zyberg):

Yes, just noticed. 6k+4 = 2(3k + 2) and that means 6 and n would have divisor 2.

ganeshie8 (ganeshie8):

Good. Let's move to step2

ganeshie8 (ganeshie8):

Before that notice this In modulus 6, the remainders 5 and -1 are sane

OpenStudy (zyberg):

Yes, I know that.

ganeshie8 (ganeshie8):

So we can say that the only possible remainders when n is divided by 6 are +-1

OpenStudy (zyberg):

Yes.

OpenStudy (zyberg):

So, what would be the second step? ;)

ganeshie8 (ganeshie8):

Step2 Let \(n = 6k \pm 1\) Then we need to show that \(( 6k \pm 1)^{2}-1\) is divisible by 24

OpenStudy (zyberg):

36k^2 +/- 12k + 1 - 1 = 24q 36k^2 + / - 12k = 24q 3k^2 + / - k = 2q

OpenStudy (zyberg):

Oh, did one step too far. But how would 36k^2 + / - 12k prove 24q?

OpenStudy (zyberg):

12(3k^2 + / - k) = 24q 3k^2 + / - k = 2q Still need to prove that left side is divisible by 2.

ganeshie8 (ganeshie8):

Looks great! Factor 3k^2 + / - k Then stare at it

OpenStudy (zyberg):

Hm... k(3k + / - 1) seems similar to n = 6k +/ - 1; n + / - 1 = 6k; k((n + / - 1)/2 + / - 1) Is that right?

ganeshie8 (ganeshie8):

\(3k^2 \pm k = 2k^2 + k^2 + k = 2k^2 + k(k+ 1) \)

OpenStudy (zyberg):

How comes that after the equal sign the k got rid of + / -?

ganeshie8 (ganeshie8):

Notice that \(2k^2\) is divisible by 2 What can you say about the expression \(k(k+1)\) ?

ganeshie8 (ganeshie8):

Oh sorry, that was a typo. Just put that - sign also

OpenStudy (zyberg):

Its going to be even! Consecutive numbers multiplied by each other give out even product. Oh my! Even + even = even and even is divisible by 2.

ganeshie8 (ganeshie8):

That'll do

OpenStudy (zyberg):

Amazing! I think that I would never get to the solution by myself ;) Thank you VERY much! :)

ganeshie8 (ganeshie8):

Np, yw :)